Find the eqaution of each line. Write the eqaution in standard form unless indicated otherwise.
Through (1,6) and (5,2); use function notation. Thanks

Answer by Alan3354(69443)

Find the eqaution of each line. Write the eqaution in standard form unless indicated otherwise.
Through (1,6) and (5,2); use function notation.
--------------------
First find the slope, m
m = diffy/diffx
m = (5-1)/(2-6)
m = -1
------
Then use either point in
y-y1 = m*(x-x1)
y-6 = -1(x-1) = -x+1
y = -x + 7 (slope-intercept form)
x+y = 7 (standard form)